

********************************************************************
UARS WEEKLY STATUS REPORT - 02 FEB 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM performed its daily solar scans of the sun during periods of
predicted UARS telemetry.
With much regret, we announce the retirement from NRL of Dr. Dianne Prinz,
SUSIM Principal Investigator. We extend to her all of our best wishes.
She will be missed. Dr. John Cook of NRL has now become SUSIM PI.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Initial tests of telemetry transfers through
PEM2K:: have been successful; however, the source of these failures has been
traced to quota limits and these have been corrected. PEM Telemetry is runing
in duplicate streams through PEM:: and PEM2K::. We find that files which
transfer from the CDHF to PEM:: and PEM2K:: are now identical.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the Monthly Status Review meeting for
UARS, held on 31 January 2001.
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 15 March 2001.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 January 2001 (Orbit 51166, GMT Day 020) through 26
January 2001 (Orbit 51270, GMT Day 026).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-None this report period.
The beta angle ranged between the angles of 55.8 degrees to 76.1 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 77 degrees on 21 January 2001 (DOY 021).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.3 and 0.9 degrees C. Additional battery performance data for each
battery on 12 January 2001 and 19 January 2001 are provided in Appendix D.
The clock error ranged from -11.2 to 10.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 02
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 07
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Five (5) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/025 show the observatory data loss to be 300 hours, 13
minutes, 18.776 seconds. This shows an overall increase of 4 minutes and 37.520
seconds since the last report period This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The decrease of 4 minutes and 37.520 seconds occurred over one (1) day:
-4 minutes and 37.520 seconds on Day 023 due to a Late AOS
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS finished flying through the high beta period on 25 January 2001 at
approximately 23:00 UTC. The first sun sensor reading showed a 650 arc second
residual, which was only a 200-250 arc second increase from the TAM residuals
received before the loss of sun. The FOT was pleased in the TAMs performance
over the high beta period and UARS attitude control was maintained. More
flight software patched are expected over the next few weeks to upgrade the TAM
to a 7th order model and to correct a slight time calculation error.
MEETINGS
The FOT attended The UARS SERB Meeting on 24 January 2001.
The FOT attended a Flight Software Meeting on 25 January 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 09 February 2001 (DOY 040). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 January 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51183 021/0443 01-051 OBC Clk Adj 20.6 to 41.2 msec/dy
51208 022/1927 01-052 TAM MAGMOD UPDATE
51225 023/2328 01-053 OBC Clk Adj 41.2 to 20.6 msec/dy
51251 025/1652 01-054 CLK Clk Adj 20.6 to 41.2 msec/dy
51254 025/2129 01-055 NBTR NOOPS D026
51261 026/0914 01-055 NBTR NOOPS D026
51266 026/1751 01-056 OBC Clk Adj 41.2 to 20.6 msec/dy
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 January 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
NBTUTEMP2 R-LO NBTR TU-2 TEMP 39 51171
QTZEPSIF1 Y-LO ZEPS Interface Temp 1 51243-4
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 January 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51183 021/0443 Clk Adj 20.6 to 41.2 msec/dy
51225 023/2336 Clk Adj 41.2 to 20.6 msec/dy
51251 025/1709 Clk Adj 20.6 to 41.2 msec/dy
51266 026/1753 Clk Adj 41.2 to 20.6 msec/dy
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 January 2001
19 Jan 2001 - GMT Day 019 Beta = 76.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ------ ------- -----
1 +728.0/+728.0 -2.42/-3.68 +0.0/+0.0 14.2 0.000 0.0
2 +11.2/-22.4 +3.23/+1.66 +15.2/-16.0 27.4 1.050 14.1
3 +11.2/-28.0 +2.29/+1.34 +14.4/-15.6 27.4 1.040 13.8
26 Jan 2001 - GMT Day 026 Beta = 58.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------ ------- -----
1 +212.8/+112.0 -3.05/-4.62 +0.0/+0.0 13.3 0.000 0.0
2 +11.2/-28.0 +4.19/+2.92 +24.0/-14.4 26.2 1.040 21.0
3 +16.8/-33.6 +3.23/+1.97 +23.2/-14.0 26.2 1.040 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 January 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-004 FOT 51220 023/1606 23512 Late Acquisition/4:02 non-rec
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
00-056 FOT 50584 347/0502 23302 Premature LOS
01-002 FOT see list Jan 01 14170 Generic Late Acquisition
01-003 FOT 51111 016/0856 N/A AP Crashed/disk crash
01-004 FOT 51220 023/1606 23512 Late Acquisition/4:02 non-rec
SEVNTFW01/04